Step by step
The question
Question: x can be 0, 1, 2, 3, with probabilities 0.25, 0.25, 0.25, a missing one: find the missing probability.
x: 0, 1, 2, 3; P(x): 0.25, 0.25, 0.25, ?
1 All P(x) add to 1
x always takes one of its values.
So all the probabilities add up to 1.
a probability table always adds up to 1
0.25 plus 0.25 plus 0.25 plus P of 3 is 1.
0.25 + 0.25 + 0.25 + P(3) = 1
The known ones add up to 0.75.
0.75 + P(3) = 1
To remove the plus 0.75, we take away 0.75 on both sides.
1 minus 0.75 is 0.25.
P(3) = 1 − 0.75 = 0.25
Careful!
The missing probability is what is left over from 1.
Careful: the missing P is what is left of 1
2 Check
Let's check: 0.25 plus 0.25 plus 0.25 plus 0.25 is 1.
So it is right!
0.25 + 0.25 + 0.25 + 0.25 = 1
Let's look back.
First: all P(x) add to 1.
Last: check.
Answer
Careful: the missing P is what is left of 1